# What Lies Beneath? Exploring the Impact of Underlying AI Model Updates in AI-Infused Systems

### v2

@@ 5. Discussion
5.1. Contextual Complexity in End-User Evaluation of AI Models
Despite the new AI model being more accurate, we did not observe unanimous support for it from the study participants. While some users clearly favored the newer model for its precision and fewer false positives, others preferred the older model due to its broader range of choices. This divergence highlights the complexity of user preferences when interacting with evolving AI models. Although there may be a possibility of neophilia (i.e., a bias towards new things or innovations) (Wikipedia contributors, 2022), our findings suggest that users were evaluating the models based on context-specific needs. Similar to other algorithm-driven platforms (Register et al., 2023; Milton et al., 2023; DeVito et al., 2018), participants crafted multiple, often divergent folk theories (see Table 3). These theories reflected attempts to understand the system’s inner workings, with users assessing the models based on how well they fit their specific use cases rather than reacting purely to novelty.
Users considered a wide array of factors in their evaluations, many of which shape the overall user experience and extend beyond traditional performance benchmarks. Factors such as cognitive load, search result volume, and the time required to sift through results were repeatedly mentioned, highlighting the importance of conducting exploratory evaluations that incorporate the actual user base. Standard technical metrics may not fully capture user experience, especially given significant individual differences. For example, participants who valued precision tended to favor the new model for its focused, high-accuracy results, whereas those who preferred a broader set of options gravitated toward the older model. To better capture these individual preferences, future research could benefit from larger-scale experiments tracking behavioral metrics, such as scrolling patterns and time spent evaluating results, and correlating these with user actions and preferences.
In both studies, participants quickly identified model shortcomings, but their evaluations were nuanced. In the first study, where users were not informed which model was in use, they struggled to differentiate between them. Even in the diary study, where model changes were explicit, inconsistent folk theories emerged, with no clear preference for one model. This contrast suggests that end-user evaluation of AI models is complex, influenced by individual needs and task context. Future work could explore long-term behavioral patterns to determine whether users continue toggling between models or eventually settle on a preferred one.
5.2. Deploying AI Model Updates in a Human-AI Collaborative Environment
Our findings underscored the importance of effectively communicating AI model updates in human-AI collaborative environments, especially when users interact with the outputs of the models on a regular basis. Across both studies, we saw that users place significant emphasis on their personal experiences, developing diverse and sometimes inconsistent folk theories about model behavior. These folk theories often diverge from standard benchmarking metrics, suggesting that conventional performance metrics alone may be insufficient to convey a model’s capabilities.
A more user-centered approach to deploying updates is necessary. One effective strategy, as evidenced in our second study, is to provide a ”try-it-on” or ”beta testing” interface that enables users to directly compare different models and explore their strengths and weaknesses. This approach allows users to identify changes firsthand and form more informed judgments. However, given the inconsistencies in folk theories, developers could also integrate social components that facilitate knowledge sharing. Enabling users to share insights and discuss their experiences could foster a more comprehensive understanding of model behavior and updates, leveraging the collective experiences of the user community, drawing social transparency ideas in AI systems (Ehsan et al., 2021).
A novel direction could involve developing platform-specific LLM agents capable of generating logs and reports of user interactions with the AI models. Aggregating these reports over time and across users could yield valuable insights into how models perform in different contexts, ultimately improving the deployment and adaptation process for both users and developers.
Finally, providing educational interfaces to help users understand key distinctions between different types of models (Kusuma et al., 2022)—such as deterministic versus stochastic behavior, which emerged in Study 1—can be crucial. This could also help users, who may not be AI-literate, to figure out key strategies on how to interact with different AI models. By combining exploratory interfaces, social learning opportunities, dynamic feedback mechanisms, and targeted education, developers can create a more supportive environment for users to effectively engage with and understand AI model updates.
@@ 6. Conclusion
Appendix A Generative AI Usage
@@
Appendix B Benchmarking the Old and New Facial Recognition Models
@@
Table 5. Model Benchmarking: Ranks and confidence scores of correct matches retrieved by Model 1 and Model 4.
Appendix C Mixed Effects Model Outputs
C.1. Distinguishing Models: Perceived Accuracy vs. Objective Characteristics
We include the detailed output of the generalized linear mixed effects model used to examine whether participants relied on observable model characteristics (e.g., latency, result count) or their own perceived accuracy to distinguish between models. This analysis corresponds to Section 3.9.2 in the main text.
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: user_model_comparison_binary ~ latency_diff_scaled + result_count_diff_scaled + perceived_accuracy_diff_scaled + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
2159.2 2186.5 -1074.6 2149.2 1752
Scaled residuals:
Min 1Q Median 3Q Max
-1.7272 -0.7875 -0.3763 0.8605 6.2023
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.4657 0.6825
Number of obs: 1757, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) -0.47974 0.07086 -6.770 1.29e-11 ***
latency_diff_scaled 0.06054 0.06022 1.005 0.315
result_count_diff_scaled 0.03753 0.05908 0.635 0.525
perceived_accuracy_diff_scaled -0.89630 0.07394 -12.122 < 2e-16 ***
[…]
Correlation of Fixed Effects:
(Intr) ltnc__ rsl___
ltncy_dff_s -0.005
rslt_cnt_d_ 0.003 -0.321
prcvd_ccr__ 0.222 -0.027 0.059
C.2. Factors influencing Perceived Accuracy
We include the detailed output of the generalized linear mixed effects model used to examine which factors influenced the perceived accuracy of a given model. This analysis corresponds to Section 3.9.4 in the main text.
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: perceived_accuracy ~ facial_match_count + rep_match_count + ns_count + diff_count + search_results_count + model_type + overall_timer +
seen_tiles + mean_target_comparison_time + latency + (1 | session_id) + (1 | task_id)
Data: df
REML criterion at convergence: 20852.3
Scaled residuals:
Min 1Q Median 3Q Max
-5.2073 -0.4940 0.1214 0.6061 5.3309
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 555.61 23.571
task_id (Intercept) 24.96 4.996
Residual 1554.75 39.430
Number of obs: 2016, groups: session_id, 252; task_id, 8
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 19.1127 2.7291 21.9115 7.003 5.10e-07 ***
facial_match_count 23.5355 1.0502 1998.9006 22.411 < 2e-16 ***
rep_match_count 16.6758 1.0244 1984.1011 16.279 < 2e-16 ***
ns_count 2.9388 1.0867 2000.0761 2.704 0.006902 **
diff_count -7.0787 1.5766 1906.5875 -4.490 7.55e-06 ***
search_results_count 5.9723 1.2112 1884.9401 4.931 8.91e-07 ***
model_typerecognition_04 14.2293 2.3095 1842.6505 6.161 8.84e-10 ***
overall_timer -6.2852 2.0136 1505.9057 -3.121 0.001834 **
seen_tiles -4.7527 1.2930 1976.4139 -3.676 0.000244 ***
mean_target_comparison_time 9.6421 1.4457 1931.4743 6.669 3.34e-11 ***
latency 0.3769 1.0954 2000.2490 0.344 0.730837
[…]
Correlation of Fixed Effects:
(Intr) fcl_m_ rp_mt_ ns_cnt dff_cn srch__ md__04 ovrll_ sn_tls mn_t__
fcl_mtch_cn 0.051
rp_mtch_cnt 0.035 0.074
ns_count 0.036 0.056 0.106
diff_count 0.015 0.121 -0.042 0.006
srch_rslts_ -0.223 -0.070 -0.072 -0.099 0.004
mdl_typr_04 -0.426 -0.119 -0.083 -0.084 -0.036 0.524
overall_tmr -0.007 -0.183 -0.144 -0.262 -0.647 0.042 0.016
seen_tiles -0.073 0.025 0.040 0.010 -0.139 -0.097 0.172 -0.335
mn_trgt_cm_ 0.019 0.119 0.096 0.087 0.414 -0.072 -0.044 -0.552 0.195
latency -0.002 0.010 0.024 0.045 0.110 -0.313 0.006 -0.162 0.008 0.070
C.3. Factors influencing model preference
We include the detailed output of the logistic regression model used to analyze the factors influencing participants’ preferences between Model A-UP and Model B-UP. The analysis evaluates the effects of trial count, match findings, and perceived accuracy on model preference, discussed in Section 3.9.5. The significant predictors are highlighted in the output below.
glm(formula = preferred_model_binary ~ model_A_trial_count +
model_A_trial_count_match_found + model_A_avg_perceived_accuracy +
model_B_trial_count_match_found + model_B_avg_perceived_accuracy,
family = binomial(link = "logit"), data = df)
Coefficients:
Estimate Std. Error z value Pr(>|z|)
(Intercept) 0.400568 1.327662 0.302 0.763
model_A_trial_count -0.308623 0.299720 -1.030 0.303
model_A_trial_count_match_found 0.075469 0.189530 0.398 0.690
model_A_avg_perceived_accuracy -0.033977 0.006874 -4.942 7.71e-07 ***
model_B_trial_count_match_found 0.092030 0.190740 0.482 0.629
model_B_avg_perceived_accuracy 0.037990 0.006681 5.686 1.30e-08 ***
[…]
(Dispersion parameter for binomial family taken to be 1)
Null deviance: 321.43 on 240 degrees of freedom
Residual deviance: 225.78 on 235 degrees of freedom
(11 observations deleted due to missingness)
AIC: 237.78
Number of Fisher Scoring iterations: 5
C.4. Impact of the underlying model on the comparison decisions
We include the detailed output of the generalized linear mixed effects model used to examine the impact of the underlying model on the comparison decisions. This analysis corresponds to Section 3.9.7 in the main text.
These are the outputs for the three models: positive vs. rest, negative vs. rest, and uncertain vs. rest.
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: positive_vs_rest ~ model_type + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
11472.1 11493.9 -5733.0 11466.1 10643
Scaled residuals:
Min 1Q Median 3Q Max
-2.1231 -0.6113 -0.4218 0.7908 5.4912
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.8151 0.9028
Number of obs: 10646, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) -1.01287 0.06717 -15.080 < 2e-16 ***
model_typerecognition_04 0.38091 0.04906 7.764 8.21e-15 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.378
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: negative_vs_rest ~ model_type + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
13100.2 13122.0 -6547.1 13094.2 10643
Scaled residuals:
Min 1Q Median 3Q Max
-3.6717 -0.8580 0.4317 0.7729 3.0907
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.8162 0.9034
Number of obs: 10646, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) 0.18270 0.06562 2.784 0.00537 **
model_typerecognition_04 -0.36777 0.04533 -8.113 4.92e-16 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.338
Generalized linear mixed model fit by maximum likelihood (Laplace Approximation) [’glmerMod’]
Family: binomial ( logit )
Formula: uncertain_vs_rest ~ model_type + (1 | session_id)
Data: df
AIC BIC logLik deviance df.resid
8750.3 8772.1 -4372.1 8744.3 10643
Scaled residuals:
Min 1Q Median 3Q Max
-0.9543 -0.4570 -0.3581 -0.2578 4.9061
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 0.5016 0.7082
Number of obs: 10646, groups: session_id, 252
Fixed effects:
Estimate Std. Error z value Pr(>|z|)
(Intercept) -1.81609 0.06201 -29.287 <2e-16 ***
model_typerecognition_04 0.08289 0.05790 1.432 0.152
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.465
C.5. Impact of Model Type on Search Behavior
We include the detailed output of the linear mixed effects models used to examine the impact of the underlying model on the search behavior, specifically on the number of comparisons made, the time spent on the search results page, and the number of results checked. This analysis corresponds to Section 3.9.7 in the main text.
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: total_num_comparisons ~ model_type + (1 | session_id)
Data: df
REML criterion at convergence: 14139.9
Scaled residuals:
Min 1Q Median 3Q Max
-2.0686 -0.1943 -0.0892 -0.0029 20.8738
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 8.64 2.939
Residual 59.12 7.689
Number of obs: 2016, groups: session_id, 252
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 5.7866 0.3079 518.8372 18.791 < 2e-16 ***
model_typerecognition_04 -1.0048 0.3509 1896.8430 -2.863 0.00424 **
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.574
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: overall_timer ~ model_type + (1 | session_id)
Data: df
REML criterion at convergence: 23280.3
Scaled residuals:
Min 1Q Median 3Q Max
-2.7329 -0.4519 -0.1463 0.2863 18.4644
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 1510 38.86
Residual 5248 72.44
Number of obs: 2016, groups: session_id, 252
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 125.079 3.378 427.941 37.032 < 2e-16 ***
model_typerecognition_04 -16.744 3.331 1856.454 -5.027 5.46e-07 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.497
Linear mixed model fit by REML. t-tests use Satterthwaites method [’lmerModLmerTest’]
Formula: seen_tiles ~ model_type + (1 | session_id)
Data: df
REML criterion at convergence: 17784.4
Scaled residuals:
Min 1Q Median 3Q Max
-1.9846 -0.4864 -0.0937 0.1384 13.1841
Random effects:
Groups Name Variance Std.Dev.
session_id (Intercept) 53.69 7.327
Residual 360.65 18.991
Number of obs: 2016, groups: session_id, 252
Fixed effects:
Estimate Std. Error df t value Pr(>|t|)
(Intercept) 21.0858 0.7632 516.8290 27.63 <2e-16 ***
model_typerecognition_04 -11.5951 0.8669 1895.9167 -13.38 <2e-16 ***
[…]
Correlation of Fixed Effects:
mdl_typr_04 -0.572

### v5

@@ 4.3.4. The trade-off between quantity and quality influenced user preferences for each model
4.4. Summary of Findings
In the diary study, participants, who had prior experience with the older modelon CWPS, were given the option to toggle between the old and new models to compare their experience. The findings revealed that while most participants preferred the new model for its higher precision and relevance, some valued the larger result set of the old model for enabling more exploratory opportunities. Perceived accuracy strongly influenced model preferences, with participants often citing facial feature alignment as a key factor for favoring the new model. However, the trade-off between quantity and quality emerged as a recurring theme, as some participants preferred the broader range of results from the old model despite its occasional inaccuracies. Interestingly, participants developed nuanced but inconsistent folk theories about model behavior, attributing differences in performance to aspects like handling of specific facial features or image quality.
@@ 5. Discussion
5.1. Model Updates in Facial Recognition Systems
Facial recognition systems have found critical applications across diverse domains such as identity verification in gig work (Watkins, 2023), law enforcement (Garvie, 2022; Harwell, 2019), and historical research (Mohanty et al., 2019b; Schwartz, 2022), where they support high-stakes decision-making and enhance operational efficiency. However, despite advancements in auditing and model design (Raji and Buolamwini, 2019, 2022), major providers like AWS and Azure often release updates without detailing their impact on downstream tasks (Azure, [n. d.]; Services, [n. d.]), leaving end-users to interpret changes independently. Human-AI teaming in facial recognition tasks relies on the complementary strengths of precise algorithmic outputs and human contextual judgment, but prior work and our findings reveal that incomplete user understanding of model behavior and reliance on subjective strategies often limit collaborative performance (Towler et al., 2023; Howard et al., 2020; Carragher et al., 2024), highlighting the importance of explicit communication and alignment in such systems.
Across both studies, participants struggled to effectively assess and distinguish between facial recognition models, highlighting challenges in user understanding of model updates. In Study 1, users relied heavily on subjective perceptions of accuracy to form their assessments, which did not translate into performance gains despite the newer model’s improved precision and recall. This reliance on personal experiences persisted in Study 2, where users toggling between explicitly labeled ”old” and ”new” models developed divergent folk theories about their properties. These findings underscore a critical limitation of silent or minimally communicated updates: without clear guidance, users may fail to align their strategies with model improvements, leaving potential performance benefits unrealized. For instance, while some participants in Study 1 spent additional time deliberating on specific matches with the newer model, this behavioral shift did not consistently enhance task outcomes. For developers, this underscores the necessity of pairing technical advancements with user-centric communication strategies, ensuring updates empower users to fully leverage improved system capabilities.
For high-stakes applications, addressing inconsistent assumptions requires targeted strategies. Educational interfaces (Kusuma et al., 2022) could address common misunderstandings about model behavior, such as non-determinism and variability. Developers should also prioritize diverse, context-specific datasets (e.g., race, gender, etc.) and move beyond static benchmarking. By combining these datasets with user simulations, streamlined evaluation workflows could produce public reports detailing both model benchmarks and simulated task outcomes, enabling informed adoption and deployment of new models.
5.2. Communicating Model Updates in AI-Infused Systems
Communicating model updates is widely recognized as a critical design challenge in AI-infused systems, yet operationalizing this in real-world settings remains underexplored. While human-AI interaction guidelines emphasize transparency in communicating changes (Amershi et al., 2019), they provide little guidance on what specific information to share — such as performance metrics, new capabilities, or expected user impacts — or how to present it effectively. Our studies underscored this challenge, as participants struggled to infer changes between models and showed no significant improvement in task performance. These findings highlight a gap in how updates are communicated, suggesting the need for more granular and targeted strategies that extend beyond simple labels like ”new” or ”old.”
Future work could explore which aspects of model updates most effectively resonate with users. Rather than focusing solely on abstract improvements (e.g., “better accuracy”), updates could be framed in terms of tangible impacts on user experiences, such as faster decision-making or enhanced time spent evaluating AI outputs. Personalized framing strategies, such as categorizing models as ”balanced,” ”creative,” or ”precise,” might also help guide user expectations and foster more intentional interactions, similar to Microsoft’s Copilot modes (Bersano, 2024). Scenarios where a ”new” model performs worse but is framed as an improvement could demonstrate novelty bias (Luo et al., 2023; Wikipedia contributors, 2022), leading users to favor it despite poorer performance. Testing how framing influences trust and skepticism remains critical.
Beyond static communication strategies, dynamic approaches informed by user behavior and perceptions hold significant promise. Vision-language model (VLM)-based agents, leveraging XAI techniques like counterfactual explanations (“what-if” scenarios) and heatmaps, could validate or challenge user-generated folk theories by linking assumptions to the model’s actual logic. For example, counterfactuals could show how alternative queries or results refute or support a folk theory, while heatmaps and concept-based explanations visually map user assumptions to system behavior (Kim et al., 2023). Large language model (LLM) agents could analyze interaction patterns to identify shallow evaluations or deviations from standard interaction patterns and adaptively intervene with tailored guidance. These dynamic systems could align user strategies with model capabilities, fostering deeper understanding and more effective human-AI collaboration.
5.3. Generalizing AI Model Update Challenges Across Domains
Our findings, while situated in the context of facial recognition, reveal broader implications for AI-infused systems where silent or subtle model updates are common. Unlike software updates, which often prompt resistance due to visible disruptions in workflows or undesired changes (Morreale and Eriksson, 2020; Rula et al., 2020), AI model updates in black-box systems present unique challenges. Participants in our study struggled to recognize changes, highlighting how opacity exacerbates the difficulty of adapting to improvements. These challenges become even more acute in high-stakes scenarios, where trust, performance, and safety are paramount.
Similar to other algorithm-driven platforms (Register et al., 2023; Milton et al., 2023; DeVito et al., 2018), participants crafted multiple, often divergent folk theories to make sense of the system’s behavior (see Table 3). These theories, while sometimes helpful, also led to incorrect assumptions about model behavior, further complicating user adaptation. In high-stakes contexts such as healthcare or autonomous vehicles, reliance on inaccurate folk theories could have severe consequences, from misinterpretation of diagnostic outputs to unsafe decision-making during critical tasks.
For example, in September 2023, Tesla’s rollout of an updated Full Self-Driving (FSD) software caused confusion and safety concerns, with drivers reporting unexpected behaviors and unclear changes (Cao, 2023). Similarly, in healthcare, hallucinations in AI transcription tools have introduced inaccuracies in patient records, forcing practitioners to adopt heightened vigilance to avoid potential errors (Edwards, 2024). These cases highlight that even silent improvements in AI accuracy may require users to adapt proactively, emphasizing the need for alignment between system behavior and user workflows.
Consistent with prior human-AI teaming work (Chattopadhyay et al., 2017; Bansal et al., 2019b), improved models did not always translate into better outcomes, as users failed to adapt their strategies to take advantage of system improvements. This misalignment underscores a common challenge across domains where users rely on preexisting behaviors rather than adjusting to updated AI capabilities. While prior work (Wang and Yin, 2023) demonstrated that users could detect explanation changes during model updates and adjust accordingly, our findings highlight the unique challenges of black-box systems. In the absence of accompanying explanations, users were unable to recognize updates, suggesting that generalizing successful strategies from explainable AI contexts may require significant adaptation for opaque systems.
In our studies, we observed a contrast between Prolific participants, who focused primarily on AI outputs, and CWPS participants, who integrated contextual reasoning, such as uniforms and locations, alongside AI assistance. Despite their expertise and long-term exposure to the older model, CWPS participants still developed inconsistent folk theories about the updated model’s behavior, illustrating the diverse beliefs users form when interpreting AI systems. For experienced users like those on CWPS, decision-making speed was not a meaningful factor, as their task emphasized thorough analysis over quick judgments. While more work is needed to determine which groups can reliably judge model changes, our findings highlight that good communication is essential for all users, irrespective of expertise, and must be tailored to diverse behaviors and beliefs to better align expectations with AI system updates.
Our work highlights the importance of granular communication in AI-infused systems, particularly for model updates, where simply notifying users of changes is insufficient. While our metrics focus on a visual retrieval task, they underscore broader challenges in AI-assisted decision-making, emphasizing the need to tailor strategies to align user behaviors with evolving system capabilities.
6. Limitations and Future Work
In Study 1, we aimed to evaluate participants’ ability to distinguish between models within a controlled environment. However, the study’s duration and design may have limited participants’ familiarity with the models, as no improvement trends were observed over the seven trials. Prolonged interactions, as seen in real-world applications, could enhance users’ ability to detect model changes over time. The design also presented trial-specific details, such as perceived accuracy ratings, which may have encouraged post-hoc justifications rather than reflecting real-time decision-making. A think-aloud study could mitigate this by capturing participants’ reasoning during interactions. While we conducted pre-planned statistical tests for recall, precision, and false positive rates (see Section 3.9.6), corrections for multiple comparisons were not applied, though the pre-planned nature and reporting of non-significant results ensured transparency. The study was also not pre-registered, which we acknowledge as a limitation. Furthermore, the lack of pre- and post-study assessments of AI literacy limits our understanding of shifts in participants’ mental models.
The controlled setting may limit ecological validity, as real-world interactions involve prolonged usage, distractions, and higher stakes. Additionally, the participant sample, recruited online, may not represent domain-specific users, such as historians, and the tasks differed from real-world workflows, limiting generalizability. The significant difference in result counts between models, reflecting their native configurations, underscores trade-offs between result quantity and quality. Artificially constraining the older model’s results could standardize comparisons but would obscure real-world behaviors. Systems like Civil War Photo Sleuth and Google Search (Google, 2024) illustrate these trade-offs, where updates reshape result characteristics and user experiences. Future work could explore how such variations influence user perceptions and performance. While our study focused on AI-assisted decision-making in a content-based image retrieval (CBIR) setting, we also acknowledge that richer forms of human-AI teaming, such as those involving agentic, iterative, and communicative AI team members (O’Neill et al., 2022), may lead to different impacts on user dynamics when models are updated.
In Study 2, a deployment study with 10 participants provided insights into model differences but was limited by the small sample size and focus on experienced users, which restricts generalizability. A larger-scale, in-the-wild deployment could yield insights into model adoption patterns and broader behaviors. While the diary study approach captured qualitative feedback, it did not log detailed interaction patterns such as scrolling and toggling. Incorporating interaction logging in future studies could address this gap. Additionally, we did not communicate the benchmarking performance of the models to avoid bias, but future studies could explore how sharing this information impacts user perceptions and performance. Finally, while the study focused on historical photo identification, further research is needed to assess how these findings generalize to other domains or AI-infused systems.
@@ 7. Conclusion
Acknowledgements.
This research was supported by NSF IIS-165196, the Dr. Dennis G. Kafura Graduate Fellowship from the Department of Computer Science at Virginia Tech, and the Center for Human-Computer Interaction at Virginia Tech. We extend our gratitude to all study participants and reviewers who provided valuable feedback during the course of this work. We also thank Dr. Alexandre Filipowicz for early discussions and suggestions, which helped shape the study design and direction.
Appendix A Appendix
A.1. Generative AI Usage
@@
A.2. Benchmarking the Old and New Facial Recognition Models
@@
Table 5. Model Benchmarking: Ranks and confidence scores of correct matches retrieved for all photos in the dataset.